# United Minds or Isolated Agents? Exploring Coordination of LLMs under Cognitive Load Theory

### v1

@@ Abstract
Large Language Models (LLMs) exhibit a notable performance ceiling on complex, multi-faceted tasks, as they often fail to integrate diverse information or adhere to multiple constraints. We posit that such limitation arises when the demands of a task exceed the LLM’s effective cognitive load capacity. This interpretation draws a strong analogy to Cognitive Load Theory (CLT) in cognitive science, which explains similar performance boundaries in the human mind, and is further supported by emerging evidence that reveals LLMs have bounded working memory characteristics. Building upon this CLT-grounded understanding, we introduce CoThinker, a novel LLM-based multi-agent framework designed to mitigate cognitive overload and enhance collaborative problem-solving abilities. CoThinker operationalizes CLT principles by distributing intrinsic cognitive load through agent specialization and managing transactional load via structured communication and a collective working memory. We empirically validate CoThinker on complex problem-solving tasks and fabricated high cognitive load scenarios, demonstrating improvements over existing multi-agent baselines in solution quality and efficiency. Our analysis reveals characteristic interaction patterns, providing insights into the emergence of collective cognition and effective load management, thus offering a principled approach to overcoming LLM performance ceilings.
@@ 1 Introduction
The increasing prevalence and capability of Large Language Models (LLMs) are transforming diverse domains, moving beyond basic text generation towards complex reasoning and problem-solving applications (Chang et al., 2024; Zhao et al., 2024; Li et al., 2024a). Aligning these powerful models with human intent and fostering effective thinking pattern is paramount for unlocking their full potential (Shen et al., 2023). In-Context Learning (ICL) is increasingly employed for alignment, offering adaptation via prompts without parameter updates (Brown et al., 2020). In this work, we adopt a broad definition of ICL, referring to the general strategy of guiding an LLM’s behavior by providing any contextual information relevant to the task to perform the task (Lampinen et al., 2024). Compared to traditional finetuning (Song et al., 2024; Lee et al., 2023), evidence suggests both methods often operate through similar mechanisms—primarily modulating the model’s thinking style rather than altering core knowledge (Lin et al., 2024; Zhao et al., 2025; Yang et al., 2024); ICL’s parameter-free nature, and adaptability make it a widely adopted paradigm for this purpose.
@@
While ICL offers flexibility, it suffers from a notable performance ceiling when applied to multi-faceted tasks requiring integration of diverse information sources (He et al., 2024; Li et al., 2023b; Kirk et al., 2023). In such scenarios, LLM agents frequently exhibit degeneration of thought, lack of diversity, or inability to follow multiple requirements (Liang et al., 2023; Huang et al., 2023; Kamoi et al., 2024; Lu et al., 2024) when using ICL. Despite increasing empirical studies on ICL’s limitations, the root causes remain under-explored. Concurrently, recent efforts to overcome the ceiling via agent-based solutions have yielded limited success, often relying on heuristics without cognitive grounding (Liu et al., 2023; Zhang et al., 2024c).
To address the first challenge—the lack of theoretical understanding behind performance ceiling—we turn to cognitive science for explanatory insight. Similar patterns of performance degradation have long been studied in cognitive science, where complex tasks involving high element interactivity often induce (Sweller, 2011, 2003). According to Cognitive Load Theory (CLT), cognitive overload happens when working memory capacity is exceeded (Baddeley et al., 1986b). Recent work suggests LLMs also exhibit bounded working memory with human-like failure modes under overload (Zhang et al., 2024b; Gong et al., 2024). These shared characteristics allow us to draw an analogy that explains the observed performance degradation in LLM agents: The performance ceiling observed in LLM agents arises when their effective cognitive load capacity is exceeded, closely mirroring the theoretical limits described by CLT.
Building on this analogical reasoning above—that the performance ceiling observed when applying In-Context Learning (ICL) to complex tasks stems from cognitive overload—we present CoThinker, a multi-agent ICL architecture that directly operationalizes insights from CLT to enhance the effectiveness of ICL and improve reasoning capacity through structured cooperation among LLM agents. Specifically, CoThinker translates the concept of collective working memory (Kirschner et al., 2018) into a practical architecture. Just as human groups distribute cognitive demands through division of labor and shared memory structures (Wilson et al., 2004; Dunbar, 1998; Tomasello, 2009), CoThinker employs specialized agents for parallel thinking and coordinates their outputs via a shared memory mechanism. This collaborative architecture enables the LLM agents to offload and manage high element interactivity, thereby mitigating the cognitive overload experienced by individual agents. To demonstrate the effectiveness of leveraging CLT in this manner, we test CoThinker on a range of complex general problem-solving tasks and specifically fabricated high cognitive load scenarios. In sum, this paper makes the following key contributions:
  • First, we are the first to explain the performance ceiling of using ICL in LLM agents by drawing a strong analogy to Cognitive Load Theory, suggesting that these limitations stem from exceeding the LLM’s effective cognitive load capacity.
@@
  • Second, based on these theoretical insights, we design and introduce CoThinker, a novel multi-agent ICL architecture. CoThinker operationalizes CLT principles, through agent specialization, transactive memory, and communication moderator to mitigate cognitive overload and enhance complex cooperation.
  • Third, we empirically validate CoThinker on complex tasks, demonstrating its ability to surpass existing multi-agent baselines. Furthermore, our analysis uncovers characteristic interaction patterns among agents, providing insights into the emergence of collective cognition within the architecture.
@@ 2.1 Multi-Agent LLM Collaboration
The development of LLMs has catalyzed significant research into multi-agent systems (MAS) where LLMs function as collaborative agents, aiming to tackle more complex problems than single agents can alone (Guo et al., 2024; Wang et al., 2024a; Qian et al., 2025). Current approaches explore various interaction structures including multi-agent debate, where agents exchange and critique ideas (Liang et al., 2023; Lu et al., 2024; Wang et al., 2024b; Du et al., 2023), iterative reflection mechanisms, enabling agents to self-correct (Shinn et al., 2023; Madaan et al., 2023; Yao et al., 2023). Role-playing and functional specialization are also prominent, assigning distinct tasks or personas to different agents to divide labor, particularly in complex, multifaceted domains (Li et al., 2023a; Qian et al., 2023; Hong et al., 2023). Architecturally, research investigates optimal communication topologies to enhance information flow (Li et al., 2024b), the dynamic formation and adaptation of agent networks (Liu et al., 2023; Wu et al., 2023), diversity of mental set (Liu et al., 2025b), and hierarchical structures for coordination (Zhang et al., 2024a). However, while these systems demonstrate advancing capabilities, their designs often draw from intuition or focus on communication efficiencies, with less explicit grounding in cognitive theories that explain effective collaboration and the management of processing limitations (Pan et al., 2025). Specifically, the systematic integration of Cognitive Load Theory (CLT) (Sweller, 2011) remains largely underexplored in the design of LLM MAS. Our work, CoThinker, directly addresses this gap by operationalizing CLT to mitigate cognitive overload in LLMs and enhance collective problem-solving.
@@ 2.2 LLM for Human Simulation
The capacity of Large Language Models (LLMs) to exhibit human-like intelligence (Liu et al., 2025a) and emulate nuanced social behaviors (Zhou* et al., 2024) is foundational to their use as artificial agents. Research has demonstrated LLMs’ ability to simulate human decision-making (Xie et al., 2024), generate believable individual and collective behaviors in social simulations (Chuang et al., 2024a), and adopt distinct personas (Chuang et al., 2024b) Critically, these parallels extend to cognitive characteristics; recent studies suggest LLMs possess bounded working memory and exhibit failure modes under cognitive overload akin to humans (Zhang et al., 2024b; Gong et al., 2024), as discussed in our introduction. Furthermore, interactions between LLM agents can mirror social psychological phenomena (Zhang et al., 2024c; Guo et al., 2024). This confluence of human-like cognitive traits, including limitations, and social capabilities provides a strong rationale for applying principles from human cognitive science—particularly theories like Cognitive Load Theory (CLT) that address cognitive limits—to the design of more effective LLM-based collaborative systems.
@@ 3 Cognitive Foundations for Enhanced LLM Performance
This section establishes the theoretical basis for our approach by drawing parallels between human cognitive limitations and observed performance ceilings in LLMs. We begin (Section 3.1) by discussing analogous constraints in working memory between humans and LLMs, a concept central to Cognitive Load Theory (CLT). Building on this, we then (Section 3.2) use CLT to interpret LLM performance degradation under complex task demands. Subsequently (Section 3.3), we examine how humans overcome individual cognitive limitations by naturally forming collective cognitive systems, and finally, we posit that these principles can inform the design of a more capable LLM architecture.
Figure 1: Analogical reasoning on how to mirror Cognitive load in human to LLM Agent to explain the performance ceiling observed when applying In-Context Learning (ICL) to LLM Agents for complex tasks, and use Cognitive Load Theory (CLT) to resolve it.
3.1 Working Memory Analogies
Human cognition relies fundamentally on working memory, a capacity-limited cognitive system associated with the prefrontal cortex, essential for temporarily holding and manipulating information during complex cognitive tasks such as reasoning and learning (Baddeley et al., 1986a; Cowan, 2010). Human working memory can only hold a limited number of information chunks simultaneously, typically around 4 to 7 (Miller et al., 1956). This system employs selective attention to filter and prioritize information (Roussy et al., 2021). LLMs exhibit intriguing functional parallels; their core attention mechanisms perform a form of sparse, selective focus on input data (Vaswani et al., 2017). Recent studies have begun to characterize a functional "working memory" in LLMs, identifying capacity limits and failure modes under high informational demands that echo human working memory phenomena (Zhang et al., 2024b; Gong et al., 2024). Thus, a key analogy emerges: both humans and LLMs operate with limited cognitive resources for the concurrent processing of information, providing a shared foundation for understanding their processing constraints. This analogy sets the stage for applying cognitive theories developed for human reasoning to interpret performance limits in LLMs (See A.1).
3.2 Cognitive Load and Performance Limits
@@
The finite nature of working memory is central to CLT (Sweller et al., 1998; Sweller, 2011). CLT distinguishes between intrinsic load, determined by the inherent complexity and element interactivity of a task, and extraneous load, which can arise from how a task or its accompanying instructions are presented. When the combined load exceeds working memory capacity, cognitive overload ensues in humans (Paas et al., 2003; Sweller, 2011). The provided guidance, meant to help, can paradoxically hinder performance if it contributes to exceeding cognitive capacity. LLM agents demonstrate analogous performance degradation when LLM agents are tasked with complex problems and guided by In-Context Learning (ICL). This often causes agents to fail at tasks they are capable of solving. For instance, tasks requiring extensive multi-step reasoning or the integration of numerous, potentially conflicting, constraints via ICL can lead to degeneration of thought, lack of diversity, or inability to follow multiple requirements (Liang et al., 2023; Huang et al., 2023; Kamoi et al., 2024; Lu et al., 2024) (See A.2). This often causes agents to fail at tasks they, in principle, are capable of solving. Drawing upon the working memory analogies and these observed patterns, we contend that such performance ceilings when applying ICL in LLMs can be understood as a manifestation of cognitive overload, where total demands surpass their effective processing capacity. To identify ways to alleviate this overload, we next examine how humans naturally overcome similar limitations through collective cognition.
3.3 Human Collective Intelligence
To surmount individual cognitive limitations, humans exhibit a capacity for collaborative problem-solving, leading to the emergence of a collective intelligence or collective mind that is more powerful than the sum of its individual constituents (Woolley et al., 2010; Malone et al., 2010; Shteynberg et al., 2023). This is not simply an aggregation of independent efforts but results from sophisticated social-cognitive abilities, including shared intentionality, theory of mind, and nuanced communication for establishing common ground (Tomasello et al., 2005; Frith and Frith, 2005). Such collective entities effectively expand cognitive resources by distributing processing. Key aspects include the formation of a collective working memory, often through Transactive Memory Systems where knowledge and responsibilities are shared (Wegner, 1987; Kirschner et al., 2018) and individuals have meta-knowledge about who knows what (Hollingshead, 2001) so that they can rely on each other for information sharing and retrieval (Hollingshead and Brandon, 2003); the engagement in parallel thinking through a division of cognitive labor, which reduces the intrinsic load on each individual (Dunbar, 2003); and the use of organized communication to integrate diverse information and maintain a shared attentional focus (Hutchins, 1995). These spontaneously formed group structures allow humans to manage complexities that would overwhelm an individual, demonstrating a natural solution to cognitive overload.
Inspired by these human collective cognitive strategies and human-LLM cognitive similarity discussed above, the subsequent section introduces CoThinker, a multi-agent ICL architecture designed to operationalize these principles to overcome LLM performance ceilings whe using ICL.

### v2

@@ Abstract
Large Language Models (LLMs) exhibit a notable performance ceiling on complex, multi-faceted tasks, as they often fail to integrate diverse information or adhere to multiple constraints. We posit that this limitation arises when the demands of a task exceed the LLM’s effective cognitive load capacity. This interpretation draws a strong analogy to Cognitive Load Theory (CLT) in cognitive science, which explains similar performance boundaries in the human mind, and is further supported by emerging evidence that reveals LLMs have bounded working memory characteristics. Building upon this CLT-grounded understanding, we introduce CoThinker, a novel LLM-based multi-agent architecture designed to mitigate cognitive overload and enhance collaborative problem-solving abilities. CoThinker operationalizes CLT principles by distributing intrinsic cognitive load through agent specialization and managing transactional load via structured communication and a collective working memory. We empirically validate CoThinker on complex problem-solving tasks and fabricated high cognitive load scenarios, demonstrating improvements over existing multi-agent baselines in solution quality and efficiency. Our analysis reveals characteristic interaction patterns, providing insights into the emergence of collective cognition and effective load management, thus offering a principled approach to overcoming LLM performance ceilings.
@@
The increasing capability of Large Language Models (LLMs) is transforming diverse domains, moving beyond basic text generation towards complex reasoning applications (Chang et al., 2024; Zhao et al., 2024; Li et al., 2024a). Aligning these powerful models with human intent and fostering effective thinking patterns is paramount for unlocking their full potential (Shen et al., 2023). In-Context Learning (ICL) is increasingly employed for this purpose. ICL adapts models via prompt-based guidance without parameter updates (Brown et al., 2020; Lampinen et al., 2024), shaping thinking patterns similarly to finetuning while avoiding retraining (Lin et al., 2024; Zhao et al., 2025; Yang et al., 2024). Its parameter-free flexibility has made ICL a widely adopted approach.
@@
While ICL offers flexibility, it suffers from a performance ceiling when applied to multi-faceted tasks requiring the integration of diverse information sources provided in-context (He et al., 2024; Li et al., 2023b; Kirk et al., 2023). In such scenarios, when guided by extensive in-context information, LLM agents frequently exhibit degeneration of thought, lack of diversity, or an inability to follow multiple requirements (Liang et al., 2023; Huang et al., 2023; Kamoi et al., 2024; Lu et al., 2024). Despite increasing empirical studies on ICL’s limitations, the root causes remain underexplored. Concurrently, recent efforts to overcome this ceiling via agent-based solutions have yielded limited success, often relying on heuristics without a firm cognitive grounding (Liu et al., 2023; Zhang et al., 2024c).
To address this gap, we turn to cognitive science for explanatory insight. Similar patterns of performance degradation under high informational demands have long been studied through the framework of Cognitive Load Theory (CLT) (Sweller, 2011; 2003). To better understand this phenomenon in LLMs, we first conduct a pilot study (Section 3). This study provides theoretical grounding by establishing the cognitive load framework for LLM performance limits and empirically examines this analogy through measurable proxies for cognitive load effects. Specifically, following CLT, we define an agent’s Working Memory (WM) as its intrinsic, capacity-limited ability to simultaneously hold and process information active in its context, analogous to the capacity of the attention mechanism (Baddeley et al., 1986b). Correspondingly, a task’s Cognitive Load (CL) is the demand it places on an agent’s WM, largely determined by the complexity and element interactivity of the information presented via ICL. When a task’s CL exceeds the agent’s WM capacity, a state of cognitive overload occurs. To examine this WM-CL analogy in LLMs, we conduct experiments measuring attention entropy and perplexity as proxies for cognitive load effects, demonstrating that LLMs exhibit patterns consistent with CLT predictions. Furthermore, recent evidence that LLMs exhibit bounded, human-like WM characteristics (Zhang et al., 2024b; Gong et al., 2024) strengthens our core posit: The performance ceiling of ICL arises when the demands of the in-context information exceed the LLM’s effective cognitive load capacity, mirroring the theoretical limits described by CLT.
Building on the CLT-grounded understanding and the empirical evidence from our pilot study, we present CoThinker, a multi-agent ICL architecture designed to mitigate the cognitive overload imposed by complex tasks. CoThinker operationalizes CLT principles through three key functions: (i) dynamic thinking style assignment that adapts to task demands rather than fixed roles, distributing intrinsic cognitive load across specialized agents, (ii) a transactive memory system that maintains shared knowledge about agent expertise and task progress, enabling cognitive offloading and reducing redundant processing, and (iii) a communication moderator that balances cognitive similarity for efficient integration with diversity for comprehensive coverage, creating small-world network properties that minimize coordination overhead while maximizing information flow. In sum, we make the following key contributions:
  • We first explain ICL failures on multi-faceted tasks as cognitive overload by formalizing a Cognitive Load Theory-based mapping from human working memory to LLM attention and context limits, and provide empirical evidence.
@@
  • We then introduce CoThinker, a novel multi-agent ICL architecture that operationalizes Cognitive Load Theory through agent specialization, a transactive memory system, and a moderated communication protocol to mitigate cognitive overload.
  • We evaluate CoThinker’s effectiveness on complex benchmarks across diverse LLMs, demonstrating its superiority over existing baselines. Additionally, we provide a component analysis that underscores its success in managing cognitive load effectively.
@@ 2.1 Multi-Agent LLM Collaboration
The rise of LLMs has spurred research into multi-agent systems (MAS), where LLMs collaborate to tackle complex problems beyond the scope of single agents (Guo et al., 2024; Wang et al., 2024a; Qian et al., 2025). Current approaches include multi-agent debates for idea exchange and critique (Liang et al., 2023; Lu et al., 2024; Wang et al., 2024b; Du et al., 2023), iterative reflection for self-correction (Shinn et al., 2023; Madaan et al., 2023; Yao et al., 2023), and role-playing or functional specialization, where agents divide tasks in complex domains (Li et al., 2023a; Qian et al., 2023a; Hong et al., 2023). Architecturally, research explores optimal communication topologies (Li et al., 2024b), dynamic agent networks (Liu et al., 2023; Wu et al., 2023), mental set diversity (Liu et al., 2025b), and hierarchical coordination (Zhang et al., 2024a). However, these designs often rely on intuition or communication efficiency, with limited grounding in cognitive theories of collaboration and processing constraints (Pan et al., 2025). Our work, CoThinker, directly addresses this gap by operationalizing Cognitive Load Theory to enhance collective problem-solving.
@@ 2.2 LLM for Human Simulation
The capacity of LLMs to exhibit human-like intelligence (Liu et al., 2025a) and emulate nuanced social behaviors (Zhou* et al., 2024) is foundational to their use as artificial agents. Research has demonstrated LLMs’ ability to simulate human decision-making (Xie et al., 2024), generate believable individual and collective behaviors in social simulations (Chuang et al., 2024a), and adopt distinct personas (Chuang et al., 2024b) Critically, these parallels extend to cognitive characteristics; recent studies suggest LLMs possess bounded working memory and exhibit failure modes under cognitive overload akin to humans (Zhang et al., 2024b; Gong et al., 2024), as discussed in our introduction. Furthermore, interactions between LLM agents can mirror social psychological phenomena (Zhang et al., 2024c; Guo et al., 2024). This confluence of human-like cognitive traits, including limitations and social capabilities, provides a strong rationale for applying principles from human cognitive science—particularly theories like Cognitive Load Theory (CLT) that address cognitive limits—to the design of more effective LLM-based collaborative systems.
@@ 3 Cognitive Foundations for Enhanced LLM Performance
This section presents our pilot study, which establishes the theoretical foundation for our approach by linking human cognitive limitations to performance ceilings in LLMs. We introduce a cognitive load model based on working memory (WM) analogies (Section 3.1), examine this analogy empirically (Section 3.2), and demonstrate how Cognitive Load Theory (CLT) can address individual limitations and guide LLM system design (Section 3.3).
Figure 1: Cognitive Load framework: Using Cognitive Load Theory (CLT), we align human and LLM cognitive load via analogy to explain ICL performance ceilings in complex tasks and to guide methods to mitigate them.
3.1 A Cognitive Load Framework for LLM Performance Limits
As depicted in the first and second blocks of Fig. 1, we propose a model that draws parallels between human cognitive limitations and performance ceilings in LLMs.
Analogous Processing Mechanisms. Human cognition relies fundamentally on working memory, a capacity-limited cognitive system associated with the prefrontal cortex that employs selective attention to filter and prioritize information during complex cognitive tasks (Baddeley et al., 1986a; Cowan, 2010; Miller et al., 1956). LLMs exhibit intriguing functional parallels through their softmax attention mechanisms, which perform selective focus on input data (Vaswani et al., 2017), with attention heads specializing in distinct processing patterns (Voita et al., 2019). Recent studies provide direct evidence that LLMs possess human-like WM characteristics, exhibiting clear limitations on concurrent information processing with performance degrading predictably as cognitive demands increase (Zhang et al., 2024b; Gong et al., 2024). More discussion in Appendix B.1B.2.
@@ 3.1 A Cognitive Load Framework for LLM Performance Limits
Cognitive Load Theory. Building upon these working memory analogies, we apply CLT (Sweller et al., 1998; Sweller, 2011) to interpret LLM performance patterns, which are derived from human WM in cognitive science. CLT distinguishes Cognitive Load (CL) between intrinsic load (determined by task complexity and element interactivity) and extraneous load (arising from instruction presentation). When the combined load exceeds working memory capacity, cognitive overload ensues. LLM agents demonstrate analogous performance degradation when tasked with complex problems via In-Context Learning (ICL): tasks requiring extensive multi-step reasoning or integration of numerous constraints can lead to degeneration of thought, lack of diversity, or inability to follow multiple requirements (Liang et al., 2023; Huang et al., 2023; Kamoi et al., 2024; Lu et al., 2024). We contend that such performance ceilings represent cognitive overload, where total demands surpass the LLM’s effective processing capacity. Examples in Appendix B.4.
3.2 Understanding Cognitive Load and Working Memory in LLMs
We empirically examine the analogy of CL and WM in LLMs. We probe into measurable proxies for cognitive load effects by definition and examine key CLT predictions regarding task and instruction complexity effects. By definition, WM handles information processing, and cognitive load represents the attention required to handle information within WM, which determines the easiness of task completion. We identified two proxies corresponding to these key characteristics:
Attention Entropy measures the diversity of the model’s attention distribution, with higher entropy indicating more distributed attention across input tokens, suggesting the model must consider multiple aspects of the input, corresponding to higher cognitive load (Zhang et al., 2025). Perplexity measures the model’s certainty of solutions, serving as a proxy for the easiness of task completion. For the Task Complexity Effect experiment, we construct Q&A pairs from AMPS Hendrycks et al. (2021a) with 4 difficulty levels (simple to complex arithmetics), controlling input length for fair comparison. For the Instruction Complexity Effect experiment, we select Q&A pairs from FLASK Ye et al. (2023) with varying instruction complexity, measuring perplexity on answers for both hard and easy tasks. See Appendix C for details of the experimental setup, results, and discussion of these proxies.
Experiment 1: Task Complexity Effect.
Task Attention Entropy
Level 1 4.44
Level 2 4.80
Level 3 5.04
Experiment 2: Instruction Complexity Effect.
Instruction PPL (Hard) PPL (Easy)
Level 1 120.50 3.37
Level 2 88.97 3.42
Level 3 85.35 3.45
Table 1: Pilot study results: (Left) Attention entropy increases with task complexity, indicating higher cognitive load. (Right) Perplexity patterns are aligned with CLT predictions: instructions help reduce cognitive load for hard tasks but show no benefit for easy tasks.
The results provide strong empirical support for our CLT-LLM analogy. Attention entropy increases with task complexity (Table 1, left), indicating that harder tasks require the model to use more information pieces, corresponding to higher cognitive load. For perplexity (Table 1, right), hard tasks show decreasing perplexity with instruction complexity, indicating that instructions help the model focus and reduce cognitive load. Easy tasks show increasing perplexity, suggesting instructions provide no benefit and may even introduce extraneous load. This aligns with CLT predictions about cognitive load management.
3.3 Collective Intelligence Principles for Cognitive Load Management
Having empirically validated the analogy, we can now leverage CLT principles to address cognitive overload in LLMs. As depicted in the final block of Fig. 1, when humans encounter tasks that exceed individual WM capacity, we employ two strategies: external tools or collective intelligence. For complex tasks where external tools are insufficient, humans naturally form collaborative cognitive systems that exceed individual capabilities, leading to the emergence of a collective mind that is more powerful than the sum of all individuals (Woolley et al., 2010; Malone et al., 2010). CLT provides principled guidance for managing cognitive load within such collective systems, particularly addressing how the introduction of new agents or coordination mechanisms can introduce extraneous load that must be carefully balanced. Appendix B.3 provides discussion from cognitive science.
This collective intelligence effectively manages cognitive load through three core mechanisms guided by CLT principles: (i) Division of Cognitive Labor through parallel thinking, allowing individuals to focus on specialized aspects of problems, thereby reducing intrinsic load per individual (Dunbar, 2003); (ii) Collective Working Memory, often through Transactive Memory Systems where knowledge and responsibilities are distributed, enabling individuals to rely on each other for information sharing while managing the extraneous load of coordination (Wegner, 1987; Kirschner et al., 2018); and (iii) Structured Communication that efficiently integrates diverse insights through organized information flow, carefully managing extraneous load to prevent cognitive overload from coordination overhead (Hutchins, 1995). Since LLMs face analogous cognitive limitations and our pilot study demonstrates measurable cognitive load effects, systematically operationalizing these CLT principles in multi-agent LLM systems should analogously provide benefits.